OpenStudy (anonymous):

Given that a = (3,4,0) and a.(a+b)=7, what is the value of a.b?

OpenStudy (unklerhaukus):

im not sure what you mean by a=(3,4,0) is that a vector or a set?

OpenStudy (anonymous):

It's a vector

OpenStudy (amistre64):

distribute, is my thought

terenzreignz (terenzreignz):

(a+b) could be many things, my preference being a+b = <1,1,0> that means b = <-2,-3,0>

OpenStudy (amistre64):

a.a + a.b = 7 a.b = 7/(a.a)

OpenStudy (amistre64):

- a.a

terenzreignz (terenzreignz):

But @amistre64 aren't there infinitely many solutions for a+b? <1,1,0> is one of them, another would be \[\Large <0,\frac74,0>\] meaning \[\Large b = \left<-3 , -\frac{9}{4},0\right>\]

OpenStudy (amistre64):

doesnt matter what a+b actually is, the questions asks to determine a.b with whats given.

terenzreignz (terenzreignz):

oh right... lol I'm sorry, I must have misread it as a+b awkwaaard I must be getting drowsy :3

OpenStudy (amistre64):

a.(a+b) 3 , 4 , 0 3+x, 4+y, 0+z ---------------- 9+3x+16+4y a.a 3 4 0 3 4 0 ----- 9+16 a.b 3 4 0 x y z ----- 3x+4y a.(a+b) = a.a+a.b = 7 a.b = 7 - a.a = 7 - 25

OpenStudy (anonymous):

Thank You : )